OpenStudy (anonymous):

(4x^2-y^2)/ (8^2+10xy+3y^2)times(4x^2-9xy-9y^2)/(2x^2-5xy-3y^2)

OpenStudy (saifoo.khan):

Factorize them!

OpenStudy (anonymous):

thanks :)

OpenStudy (saifoo.khan):

\[\frac{(2 x-y) (2 x+y)}{(2 x+y) (4 x+3 y)} \times \frac{(4 x+3 y)(x-3 y) }{(x-3 y) (2 x+y)}\]Phew.

OpenStudy (saifoo.khan):

You missed a "x" with 8x^2.

OpenStudy (anonymous):

LOL oops thank you

OpenStudy (saifoo.khan):

I'm pretty sure you can carry on from here?

OpenStudy (anonymous):

yep thanks again

OpenStudy (saifoo.khan):

Welcome.